Concurrency: Practice and Experience Status Report — February 16, 1999

(C332) “Distributed Context Switching: A Technique to Improve Turnaround Times of Distributed Computa-

tions” Authors: Donald McLaughlin and Parth Dasgupta Current Status: Referee comments to authors (9/17/97).

(C341) “Automatic Task-Level Parallelization with pTASK”

Authors: Tarek S. Abdelrahman, and Sum Huynk Current Status: Referee comments to authors (7/3/97).

(C345) “CRAFT: A Framework for F90/HPF Compiler Optimizations”

Authors: Marina Chen, James Cowie, and Jan-Jan Wu Current Status: Referee comments to authors (7/10/97).

(C348) “Dynamic Load Balancing for Object-Based Parallel computations”

Authors: Michele Di Santo, Franco Frattolillo, Wilma Russo, and Eugenio Zimeo Current Status: Referee comments to authors (9/17/97).

(C364) “PRAM Programming: In Theory and In Practice”

Authors: D. S. Lecomber, C. J. Siniolakis, and K. R. Sujithan Current Status: Referee comments to authors (10/27/97).

(C373) “Combining Linear Data Patterns for Accessing Parallel Memories through Artitrary Multistage Net-

works” Authors: Mayez Al-Mouhamed and Lubomir Bic Current Status: Referee comments to authors (1/22/98).

(C377) “Developing Parallel Applications in a Distributed Heterogeneous Environment”

Authors: Alvaro E. Campos, Miguel Nussbaum and Ricardo Schmutzer Current Status: Referee comments to authors (7/31/98).

(C379) “A Vector-based Strategy for Dynamic Resource Allocation”

Authors: Jerrell Watts, and Stephen Taylor Current Status: Referee comments to authors (7/29/98).

(C384) “Parallel Sensitivity Analysis for DAEs with Many Parameters”

Authors: Wenjie Zhu, and Linda Petzold Current Status: Referee comments to author (9/27/98).

(C385) “The Design and Implementation of the Parallel Out-of-core ScaLAPACK LU, QR and Cholesky Fac-

torization Routines” Authors: Eduardo D’Azevedo, and Jack Dongarra Current Status: Referee comments to author (10/12/98).

(C387) “A Tool for Parallel Dynamic Programming Algorithms”

Authors: D. Gonzalez, F. Almeida, J. Roda, and C. Rodriguez Current Status: Referee comments to author (10/1/98).

(C392) “Synchronization Transformations for Parallel Computing”

Authors: Pedro C. Diniz and Martin C. Rinard Current Status: Letter to referees J. Subhlok, and R. Fowler (8/24/98). Comments received from M. Oudshoorn (2/1/99)—reject.

(C393) “After You, Alfonse and Gustave: A Toolkit for Developing Distributed Algorithms”

Authors: Steven Burdette and Tracy Camp Current Status: Letter to referees P. B. Hansen (8/13/98). Comments received from G. Thiruvathukal (9/29/98)—reject.

(C394) “Designing Shared-Memory Communication Devices for MPI Message-Passing Libraries” Authors: Boris V. Protopopov and Anthony Skjellum Current Status: Letter to referee W. Gropp (8/16/98). Comments received from V. Sunderam (10/19/98)—accept w/changes.

(C395) “Object-Oriented Analysis and Design of the Message Passing System”

Authors: A. Skjellum, D. Wooley, Z. L, M. Wolf, P. Bangalore, A. Lumsdaine, J. Squyres, and B. McCandless Current Status: Letter to referee B. Carpenter (8/16/98). Comments received from M. Snir (9/5/98)— accept w/changes.

(C396) “On the Implementation of the Opus Coordination Language”

Authors: Erwin Laure, Matthew Haines, Piyush Mehrotra, and Hans Zima Current Status: Letter to referees K. Hawick, C. Polychronopoulos, I. Foster, F. Berman, and C. Koelbel (11/25/98).

(C397) “Reproducible Execution of SR Programs”

Authors: Ronald A. Olsson Current Status: Letter to referees B. Miller, J. Mellor-Crummey, and F. Vaughan (11/25/98).

(C398) “A Load Balanced Distributed Computing System”

Authors: M. Bozyigit, J. Al-ghamdi, H. Barada, and M. Ghouseuddin Current Status: Letter to referees D. Duke and L. Turcotte (10/13/98).

(C399) “A Distributed Coordinated Atomic Action Scheme”

Authors: Alexander Romanovsky, and Avelino Zorzo Current Status: Letter to referees N. Lynch, and K. Maciunas (11/25/98).

(C400) “Distributed Control Parallelism in Multidisciplinary Aircraft Design”

Authors: Denitza T. Krasteva, Layne T. Watson, Chuck Baker, Bernard Grossman, William H. Mason, and Raphael T. Haftka Current Status: Letter to referees M. Salas, J. Dennis, and W. Furmanski (10/26/98); D. Bailey, and R. Lusk (11/16/98).

(C401) “Engineering Parallel Symbolic Programs in GpH”

Authors: Hans-Wolfgang Loidl, Philip W. Trinder, Kevin Hammond, Sahalu B. Junaidu, Richard G. Morgan, Simon L. Peyton Jones Current Status: Abstract to Board (11/3/98).

(C402) “A Concurrent Process Creation Service to Support SPMD Based Parallel Processing on COWs” Authors: M. Hobbs, and A. Goscinski Current Status: Abstract to Board (11/29/98).

(C403) “Theoretical Analysis of Optimal Striping Unit Size for Disk Array”

Authors: Hai Jin, and Kai Hwang Current Status: Letter to referees Alok Choudhary and Craig Patten (2/14/99).